# A Survey on Efficient Federated Learning Methods for Foundation Model Training

### v1

@@ 3.3 Prompt Tuning
An equivalently efficient method for tuning pre-trained models to a downstream task is prompt tuning. Here, we use binary sentiments subsequent to masked-language-modeling to achieve high-quality results Lester et al. (2021). As such, the model remains entirely frozen, and we only train for a very small number of tokens that are being prepended to each embedded input query to improve the output quality. In contrast to fine-tuning, this method does not interfere with the model architecture or parameters. Lester et al. [2021] show that the effects of prompt tuning on model performance in centralized settings become better with larger models, i.e., for large FMs, prompt tuning bears significant potential.
@@ 3.4 Instruction Tuning
All fine-tuning methods require refinement of the pre-trained Natural Language Processing (NLP) FM to work well on downstream tasks. However, for some use cases, this is necessary due to their application in fields involving highly sensitive and strictly regulated data or requiring a very high model performance.
This is where instruction tuning comes into play as a technique that uses high-quality data. For instance, GPT-4 OpenAI (2023) uses Reinforcement Learning with Human Feedback (RLHF), which provides a high level of control over the training process and helps increase model quality. This, in general, requires very precise instructions to achieve the desired results. With FedIT, Zhang et al. [2023a] study instruction tuning on LLaMA-7B in an FL application over heterogeneous client tasks, e.g., learning brainstorming and text summarization on different clients in a single system at the same time. Their results show that the additional context on a downstream task generated with federated instruction tuning provides net benefits over central training only, even in heterogeneous settings. However, it is important to note that these results were produced over a single dataset only, Dollybricks-15k.
@@ 3.5 Discussion
As more and more open-source FMs become available for unrestricted use (e.g., Alpaca, Falcon), it is unlikely that full model training will be a common use case since training an FM from scratch in a centralized setting is very challenging. Also, for very sensitive data, where privacy is a key concern, like healthcare records, providers are establishing regulatory schemes for researchers to access and work with the data already de Kok et al. (2023).
@@
Therefore, we see a priority in improving upon PEFT for downstream tasks, for instance, by introducing new and enhancing existing algorithms to remove the currently required warm-up times to improve the performance of LoRA in non-IID data environments. Computer Vision (CV) applications may benefit from exploring prompt tuning for vision transformers Jia et al. (2022). Also, we find significant challenges for instruction tuning as data quality is a general issue in FL Longpre et al. (2023) and access to trained staff, as it is required for RLHF, is hardly available in a real-world federated setting.
@@
Furthermore, PEFT also positively affects communication efficiency as it significantly reduces the number of trainable (and thus communicable) parameters. As such, we now see the computational and communication efficiency grow closer together for FL and FMs.
Table 2: Communication-efficient FL methods. Their centralized learning pendants are often tied to specific domains: CV Habib et al. (2023), NLP He et al. (2021), Audio Perez et al. (2020), and Graph data Yang et al. (2021).
Enhancement Underlying FL Max. Model Domain
@@
Algorithm Method(s) Aggregation Strategy Size (Params) CV NLP Audio Graph
Centralized Learning MP, Q, S, LRC $>100\mathrm{B}$
FedOBD Chen et al. (2023) Q Weighted Average 17M
PruneFL Jiang et al. (2022) MP Weighted Average 132M
FedPM Isik et al. (2022) MP Weighted Average 12M
FedTiny Huang et al. (2022) MP Weighted Average 132M
SoteriaFL Li et al. (2022) S FedSGD 0.05M
FjORD Horváth et al. (2021) MP Weighted Average 11M
H-FL Yang (2021) LRC Weighted Average 138M
FedPAQ Reisizadeh et al. (2020) Q FedSGD 0.09M
HeteroFL Diao et al. (2020) MP Weighted Average 11M
LotteryFL Li et al. (2020) MP Weighted Average 138M
4 Importance of Communication for FMs
@@
With FMs, models have exponentially grown from a few million to several billion parameters to be able to resemble complex tasks. For FL, this specifically means that the communication load between clients and servers has grown significantly, even though with PEFT, there is not necessarily the need to communicate an entire model. However, when fine-tuning a multi-billion parameter FM with adapter-based methods, we still need to facilitate communication for more than a million parameters. For cross-device scenarios involving more than 1,000 clients per training round, the data traffic can quickly overstrain a server’s network capacity and potentially incur significant communication costs for data transfer from the edge to a cloud Xu et al. (2023); Isenko et al. (2023). Therefore, efficient communication and training design is essential for FL systems handling FMs. We distinguish two major efficiency methods: model pruning and full model compression. A detailed overview of studies on efficient communication in FL is provided in Table 2.
@@ 4.1 Model Pruning
The objective of model pruning (MP) is to retain and communicate only parts of a DL model that are relevant to a certain task. The reduction of parameters with this technique reduces the communication effort and the computational intensity Zhu and Gupta (2017). However, the success of pruning highly depends on the underlying data. Pruning client models without coordination may deny convergence with heterogeneous non-IID data in FL systems. Jiang et al. [2022] introduce PruneFL, a two-stage procedure to realize model pruning in FL systems. The first stage is carried out on a powerful client to find a common initialization for the model and generate the importance-based pruning mask. This mask is then iteratively refined over multiple FL training rounds under the consideration of all clients. The experimental results with PruneFL show two remarkable results: (I) The models have a shorter time to accuracy over the same task with PruneFL than with FedAvg, which is attributable to the higher degree of model specialization. (II) Since the model size is reduced, one would expect additional effects on faster computation, but there is limited hardware support for sparse matrix multiplications in training. Therefore, there are no computational benefits of PruneFL to this point. However, this may change with new hardware, such as sparse Tensor cores Zhu et al. (2019). With FedTiny, Huang et al. [2022] present an approach that works identically to PruneFL, except for them swapping the common initialization procedure with using batch normalization values of clients to choose a common initialization. Furthermore, FjORD Horváth et al. (2021) and HeteroFL Diao et al. (2020) provide similar approaches to pruning.
@@
Model pruning has also been discussed extensively for FM fine-tuning outside of FL Lagunas et al. (2021); Sanh et al. (2020). As existing pruning approaches have shown strong benefits to delivering on-par model performance compared to fine-tuning the full model, this is a promising direction to combine federated PEFT with highly efficient pruning techniques to further enhance communication. Along with pruning, sparse tensor computation hardware can also lower computational loads.
@@ 4.2 Full Model Compression
Model pruning is prone to omit segments of a DL model that may become relevant at a later stage. This originates from domain shifts Peng et al. (2020) and might require preserving the full model with all its parameters. For this, there are three frequently discussed techniques for full model compression in FL systems: Quantization, Sparsification, and Low-rank Compression.
Quantization (Q). The first work towards dynamic quantization is FedPAQ Reisizadeh et al. (2020), which combines FedAvg with strong quantization guarantees, where $Q$ represents the quantization term for a local model update,
@@
While DL models often operate on full precision (32-bit), this high degree of detail is not necessarily required Zhou et al. (2018). FedPAQ leverages this to reduce the communication intensity of FL applications. $Q$ calculates the optimal float precision of a model update to preserve all required information: $Q(w)=\mathinner{\!\left\lVert w\right\rVert}\cdot\mathrm{sign}(w)\cdot\xi(w,s)$, as proposed in QSGD by Alistarh et al. [2017]. $\xi$ formulates a stochastic process to dynamically tune $s$, the level of precision. FedPAQ has a significantly lower time to accuracy than QSGD, which is attributable to dynamizing $s$. However, it must be noted that dynamic quantization only yields benefits for communication. Depending on the infrastructure, the model updates have to be cast back to full precision, creating additional computational overhead on the client and server. Also, the method has been only tested with small models ($<100\mathrm{K}$ parameters).
@@
Sparsification (S). While model pruning and sparsification technically have the same objective, pruned models do not necessarily resemble sparse models. A model is sparse once more than 50% of weights are set to 00 Frankle and Carbin (2019). However, pruning can also change the model architecture, i.e., not return the full model. Since it lends its idea from Frankle and Carbin (2019), FedPM Isik et al. (2022) (see section 4.1) can be considered as a model sparsification technique but does not necessarily lead to sparse networks and may return partial networks. SoteriaFL Li et al. (2022) guarantees sparse networks while maintaining differential privacy. Equation 7 is amended in such a way that $Q(w^{n}_{t+1}-w_{t})$ is replaced by $C(w^{n}_{t+1}+\mathcal{N}(0,\sigma^{2}\mathcal{I}))$ with $C$ resembling a sparse client update through shifted compression that has proven to improve convergence of DL models in FL settings compared to direct compression Mitra et al. (2021); Mishchenko et al. (2019). Overall, SoteriaFL mitigates the trade-off between model utility and compression, i.e., the differentially private models converge faster in stricter compression regimes than previously existing non-compressed differentially private approaches.
Low-rank Compression (LRC). H-FL Yang (2021) introduces a multi-level FL system with low-rank compression for full models by applying singular value decomposition to the feature matrix. This comes at the cost of model accuracy. To mitigate these effects, Yang [2021] includes a bias correction mechanism before each FL training round. However, it is to be noted that H-FL is a hierarchical learning framework that trains shallow models on the clients and employs mediating servers to connect the shallow client models with additional layers to form deep neural networks. The server is responsible for aggregating two models: the client models and the deep mediator network. Depending on the model architecture and nature of the task, maintaining two models could be cumbersome and incur additional communication overhead.
@@ 4.3 Discussion
To date, advancements in communication efficient methods for FL systems predominantly focus on training small full models. The communication paradigm shifts with the emergence of FMs in FL applications. With fine-tuning tasks, we only need to train a small fraction of model parameters, and thus, only trainable parameters have to be communicated. However, each trainable parameter usually contains a high degree of information for a downstream task. As such, the effectiveness of model pruning techniques is unclear as they would cut away fine-tuned parameters. At the same time, full-model compression techniques do not alter the model structure but rather reduce the parameter precision. Thus, these techniques can be used with FL applications with FMs in order to further reduce the size of the communicated updates. However, since fine-tuning is used for learning specific details in a downstream task, it is open to what degree we can compress PEFT model updates without removing essential information. Since PEFT is already sensitive to non-IID, efficient communication techniques can cause compounding effects as they remove information from trainable parameters.
Table 3: Current capabilities of state-of-the-art FL framework with respect to our taxonomy and their ability to run in resource-limited environments. The table is sorted by the frameworks’ ability to train FMs. Key: DP = Differential Privacy, HEC = Homomorphic Encryption, SMPC = Secure Multi-Party Computation.
@@
Clara DP, HEC
FedML DP, HEC PEFT
FederatedScope PEFT
Flower DP, SMPC PEFT
FATE HEC, SMPC PEFT
@@
TFF DP
@@
5 FMs in current FL frameworks
The backbones for making FL applications available to a broad audience are FL frameworks that implement recent advancements in FL research. Therefore, we investigate widely used frameworks for their FM readiness and capability with regard to computational and communication efficiency (Table 3).
Generally, all FL frameworks could handle FMs. Yet, only half implement efficient training methods. To further drive the adoption of FL in times of FMs, the frameworks need to improve both computationally and communication-efficient methods for training.
Clara is a proprietary FL framework designed specifically for medical applications and to work with NVIDIA GPUs NVIDIA (2021). The focus on powerful infrastructure makes training FMs in an FL setting possible. However, the Clara SDK does not include FL-specific optimizations to reduce computational or communication efforts specific to FMs. The final set of FL frameworks, characterized by their active open-source community, FATE Fan et al. (2023); Liu et al. (2021), FedML He et al. (2020), TensorFlow Federated (TFF) Google (2019), FederatedScope Kuang et al. (2023); Xie et al. (2023) and Flower Beutel et al. (2020), have adopted recent advancements in FL research. The frameworks especially allow for PEFT of FMs with LoRA.
Substra Galtier and Marini (2019), PySyft Ziller et al. (2021), OpenFL Foley et al. (2022), and IBM FL Ludwig et al. (2020), in their versions as of 2023, focus on training smaller FL tasks with several hundred thousand up to a few million parameters and, therefore, do not provide adapters for FM workloads with more than 100M parameters.
Yet, a consistent observation across all frameworks is their lack of efficient communication techniques. Workloads with FMs will significantly increase communication costs, and the growing use cases involving resource-constrained edge and IoT devices require high efficiency for computation and communication.

### v3

@@ 3.3 Prompt Tuning
Prompt tuning is another efficient method for tuning pre-trained models to a downstream task. Here, we use binary sentiments subsequent to masked-language-modeling to achieve high-quality results Lester et al. (2021). As such, the model remains entirely frozen, and we only tweak the prompts (a very small number of tokens) that are being prepended to each embedded input query to improve the output quality. In contrast to fine-tuning, this method does not interfere with the model architecture or parameters. Lester et al. [2021] show that the effects of prompt tuning on model performance in centralized settings become better with larger models, i.e., for large FMs, prompt tuning bears significant potential.
@@ 3.4 Instruction Tuning
Some applications work with highly sensitive and protected data or require a very high model performance. The previously mentioned fine-tuning techniques may not yield sufficient results in these cases.
This is where instruction tuning comes into play as a technique that uses high-quality data. For instance, GPT-2 Radford et al. (2019) uses Reinforcement Learning with Human Feedback (RLHF). RLHF is a multi-stage process where an FM is initially trained on supervised data. In the second step - reward model training - the FM generates outputs over a given prompt, which a user then ranks based on their preference. With this, the model learns human preferences. In the third step - proximal policy optimization - the model trains self-supervised for a maximum reward Zheng et al. (2023).
With FedIT, Zhang et al. [2023a] study instruction tuning on LLaMA-7B in an FL application over heterogeneous client tasks, e.g., learning brainstorming and text summarization on different clients in a single system at the same time. Their results show that the additional context on a downstream task generated with federated instruction tuning provides net benefits over central training only, even in heterogeneous settings. However, these results were only produced over a single dataset, Dollybricks-15k.
@@ 3.5 Discussion
As more open-source FMs become available for unrestricted use (e.g., Alpaca Taori et al. (2023), Falcon Penedo et al. (2023)), it is unlikely that full model training will be a common use case since training an FM from scratch is very challenging, even in a centralized setting.
@@
Therefore, we see a priority in improving upon PEFT for downstream tasks, for instance, by introducing new and enhancing existing algorithms to remove the currently required warm-up times to improve the performance of LoRA in non-IID data environments. Computer Vision (CV) applications may benefit from exploring prompt tuning for vision transformers Jia et al. (2022). Also, we find significant challenges for instruction tuning as data quality is a general issue in FL Longpre et al. (2023), and access to human preferences, as it is required for RLHF, is hardly available in a real-world federated setting without incentive schemes.
@@
Furthermore, PEFT also positively affects communication efficiency by significantly reducing the number of trainable (and thus communicable) parameters. As such, we now see computational and communication efficiency growing closer for FL and FMs, but not to a sufficient degree. As such, there is still a stark need to develop new approaches to use PEFT to improve computational and communication efficiency.
Enhancement Underlying FL Max. Model Communication Domain
@@
Algorithm Method(s) Aggregation Strategy Parameters Savings vs. FedAvg CV NLP Audio
Centralized Learning MP, Q, S, GP $\geq 7\mathrm{B}$ Accepted for publication at IJCAI 2024. Please find the paper in the official proceedings: https://doi.org/10.24963/ijcai.2024/919 Accepted for publication at IJCAI 2024. Please find the paper in the official proceedings: https://doi.org/10.24963/ijcai.2024/919 Accepted for publication at IJCAI 2024. Please find the paper in the official proceedings: https://doi.org/10.24963/ijcai.2024/919
FedKSeed Qin et al. (2023) GP Weighted Average 3B $\geq 99\%$ Accepted for publication at IJCAI 2024. Please find the paper in the official proceedings: https://doi.org/10.24963/ijcai.2024/919
FedOBD Chen et al. (2023) Q Weighted Average 17M $89\%$ Accepted for publication at IJCAI 2024. Please find the paper in the official proceedings: https://doi.org/10.24963/ijcai.2024/919 Accepted for publication at IJCAI 2024. Please find the paper in the official proceedings: https://doi.org/10.24963/ijcai.2024/919
PruneFL Jiang et al. (2022) MP Weighted Average 132M $80\%$ Accepted for publication at IJCAI 2024. Please find the paper in the official proceedings: https://doi.org/10.24963/ijcai.2024/919
FedPM Isik et al. (2022) MP Weighted Average 12M $98\%$ Accepted for publication at IJCAI 2024. Please find the paper in the official proceedings: https://doi.org/10.24963/ijcai.2024/919
FedTiny Huang et al. (2022) MP Weighted Average 132M $97\%$ Accepted for publication at IJCAI 2024. Please find the paper in the official proceedings: https://doi.org/10.24963/ijcai.2024/919
SoteriaFL Li et al. (2022) S FedSGD 0.05M N/A Accepted for publication at IJCAI 2024. Please find the paper in the official proceedings: https://doi.org/10.24963/ijcai.2024/919
FjORD Horváth et al. (2021) MP Weighted Average 11M $98\%$ Accepted for publication at IJCAI 2024. Please find the paper in the official proceedings: https://doi.org/10.24963/ijcai.2024/919 Accepted for publication at IJCAI 2024. Please find the paper in the official proceedings: https://doi.org/10.24963/ijcai.2024/919
FedPAQ Reisizadeh et al. (2020) Q FedSGD 0.2M $90\%$ Accepted for publication at IJCAI 2024. Please find the paper in the official proceedings: https://doi.org/10.24963/ijcai.2024/919
HeteroFL Diao et al. (2020) MP Weighted Average 11M $98\%$ Accepted for publication at IJCAI 2024. Please find the paper in the official proceedings: https://doi.org/10.24963/ijcai.2024/919
LotteryFL Li et al. (2020) MP Weighted Average 138M $50\%$ Accepted for publication at IJCAI 2024. Please find the paper in the official proceedings: https://doi.org/10.24963/ijcai.2024/919
Table 2: Communication-efficient FL methods. Their centralized learning pendants are often tied to specific domains: CV Habib et al. (2023), NLP He et al. (2021), and Audio Perez et al. (2020).
4 Communication Efficiency
@@
With FMs, models have exponentially grown from a few million to several billion parameters to be able to serve multi-modal tasks Bommasani et al. (2021). For FL, this specifically means that the communication load between clients and servers has grown significantly, even though with PEFT, there is not necessarily the need to communicate an entire model. However, when fine-tuning a billion-parameter FM with adapter-based methods, we still need to facilitate communication for millions of parameters. For cross-device scenarios involving more than 1,000 clients per training round, the data traffic can quickly overstrain a server’s network capacity and potentially incur significant communication costs for data transfer from the edge to a cloud Xu et al. (2023); Erben et al. (2023). Therefore, efficient communication and training design is vital for future FL systems. We distinguish two major efficiency methods: model pruning and full model compression. A detailed overview of studies on efficient communication in FL is provided in Table 2.
@@ 4.1 Model Pruning
The objective of model pruning (MP) is to retain and communicate only parts of a DL model that are relevant to a certain task. The reduction of parameters with this technique reduces the communication effort Zhu and Gupta (2017). However, the success of pruning highly depends on the underlying data. Pruning client models without coordination may deny convergence with heterogeneous non-IID data in FL systems. Jiang et al. [2022] introduce PruneFL, a two-stage procedure to realize model pruning in FL systems. The first stage is carried out on a powerful client to find a common initialization for the model and generate the importance-based pruning mask. This mask is then iteratively refined over multiple FL training rounds under the consideration of all clients. The experimental results with PruneFL show two remarkable results: (I) The models have a shorter time to accuracy over the same task with PruneFL than with FedAvg, which is attributable to the higher degree of model specialization. (II) Since the model size is reduced, one would expect additional effects on faster computation, but there is limited hardware support for sparse matrix multiplications in training as they are required in PruneFL. Therefore, PruneFL has no computational benefits to this point. However, this may change with new hardware, such as sparse Tensor cores that support PruneFL’s dynamic pruning approach Zhu et al. (2019). With FedTiny, Huang et al. [2022] present an approach that works identically to PruneFL, except for them swapping the common initialization procedure with using batch normalization values of clients to choose a common initialization. Furthermore, FjORD Horváth et al. (2021) and HeteroFL Diao et al. (2020) provide similar approaches to pruning.
@@
Model pruning has also been discussed extensively for FM fine-tuning outside of FL Lagunas et al. (2021); Sanh et al. (2020). As existing pruning approaches have shown strong benefits to delivering on-par model performance compared to fine-tuning the full model, this is a promising direction to combine federated PEFT with highly efficient pruning techniques to further enhance communication. Along with pruning, sparse tensor hardware can lower computational loads.
@@ 4.2 Full Model Compression
Model pruning is prone to omit segments of a DL model that may become relevant at a later stage. This originates from domain shifts Peng et al. (2020) and might require preserving the full model with all its parameters. For this, three frequently discussed techniques for full model compression in FL systems are Quantization, Sparsification, and Gradient Projection.
Quantization (Q).
The first work towards dynamic quantization is FedPAQ Reisizadeh et al. (2020), which combines FedAvg with strong quantization guarantees, where $Q$ represents the quantization term for a local model update,
@@
[…]
@@
Sparsification (S).
While model pruning and sparsification technically have the same objective, pruned models do not necessarily resemble sparse models. A model is sparse once more than 50% of weights are set to 00 Frankle and Carbin (2019). However, pruning can also change the model architecture, i.e., not return the full model. Since it lends its idea from Frankle and Carbin (2019), FedPM Isik et al. (2022) (see Section 4.1) can be considered as a model sparsification technique but does not necessarily lead to sparse networks and may return partial networks. SoteriaFL Li et al. (2022) guarantees sparse networks while maintaining differential privacy. Equation 7 is amended in such a way that $Q(w^{n}_{t+1}-w_{t})$ is replaced by $C(w^{n}_{t+1}+\mathcal{N}(0,\sigma^{2}\mathcal{I}))$ with $C$ resembling a sparse client update through shifted compression that has proven to improve convergence of DL models in FL settings compared to direct compression Mitra et al. (2021); Mishchenko et al. (2019). Overall, SoteriaFL mitigates the trade-off between model utility and compression, i.e., the differentially private models converge faster in stricter compression regimes than previously existing non-compressed differentially private approaches.
Gradient projection (GP).
Qin et al. [2023] introduce FedKSeed tailored to efficiently train FMs in FL applications. They do so by using seeds in the form of scalar vectors to create gradient projections. As only the scalar vectors have to be transmitted, the total amount of communication is reduced by $\geq 99\%$ compared to applications that would send the original multi-dimensional gradients. This is the first technique that enables communication efficient FL applications with FMs, regardless of the training regime (pre-training or fine-tuning).
@@ 4.3 Discussion
To date, advancements in communication-efficient methods for FL systems have predominantly focused on training small, full models. The communication paradigm shifts with the emergence of FMs in FL applications. With fine-tuning tasks, we only need to train a small fraction of model parameters, and thus, only trainable parameters have to be communicated. However, each trainable parameter usually contains a high degree of information for a downstream task. As such, the effectiveness of model pruning techniques is unclear as they would cut away fine-tuned parameters. An unexplored space in FL research is the pruning of FMs with subsequent fine-tuning for downstream tasks. Pruning FMs can lead to smaller transformer layers and, consequently, smaller PEFT layers with fewer parameters. In turn, this could positively affect computational and communication efficiency. Full-model compression techniques do not alter the model structure but rather reduce the parameter precision. Thus, these techniques can be used with FL applications and FMs in order to further reduce the size of the communicated updates. Furthermore, Qin et al. [2023] have shown that gradient projection is a promising direction to compress model updates without sacrificing significant information. This can be beneficial for PEFT applications as we have small specific adapters for downstream tasks. However, the remaining key challenge is the effect of non-IID data on PEFT. Potential compounding effects of communication compression and lossy compression remain open for investigation.
5 Are FL Frameworks Ready for FMs?
The backbones for making FL applications available to a broad audience are FL frameworks implementing recent advancements in FL research. We investigate widely used frameworks for their FM readiness and progress in integrating computational and communication efficiency (Table 3).
@@ 5 Are FL Frameworks Ready for FMs?
FLARE DP, HEC Currently, none of the Frameworks implements communication efficient FL methods. Accepted for publication at IJCAI 2024. Please find the paper in the official proceedings: https://doi.org/10.24963/ijcai.2024/919
FedML DP, HEC PEFT Accepted for publication at IJCAI 2024. Please find the paper in the official proceedings: https://doi.org/10.24963/ijcai.2024/919 Accepted for publication at IJCAI 2024. Please find the paper in the official proceedings: https://doi.org/10.24963/ijcai.2024/919
FederatedScope PEFT Accepted for publication at IJCAI 2024. Please find the paper in the official proceedings: https://doi.org/10.24963/ijcai.2024/919 Accepted for publication at IJCAI 2024. Please find the paper in the official proceedings: https://doi.org/10.24963/ijcai.2024/919
Flower DP, SMPC PEFT Accepted for publication at IJCAI 2024. Please find the paper in the official proceedings: https://doi.org/10.24963/ijcai.2024/919 Accepted for publication at IJCAI 2024. Please find the paper in the official proceedings: https://doi.org/10.24963/ijcai.2024/919
FATE HEC, SMPC PEFT Accepted for publication at IJCAI 2024. Please find the paper in the official proceedings: https://doi.org/10.24963/ijcai.2024/919 Accepted for publication at IJCAI 2024. Please find the paper in the official proceedings: https://doi.org/10.24963/ijcai.2024/919
@@
TFF DP Accepted for publication at IJCAI 2024. Please find the paper in the official proceedings: https://doi.org/10.24963/ijcai.2024/919
@@
Table 3: Current capabilities of state-of-the-art FL framework with respect to our taxonomy and their ability to run in resource-limited environments. Key: DP = Differential Privacy, HEC = Homomorphic Encryption, SMPC = Secure Multi-Party Computation.
In theory, all FL frameworks could handle FMs with sufficient hardware availability. Yet, only some implement efficient training methods. To further drive the adoption of FL in times of FMs, the frameworks need to improve both computationally- and communication-efficient methods for training. FL frameworks that are characterized by their active open-source community, FLARE Roth et al. (2022), FATE Fan et al. (2023); Liu et al. (2021), FedML He et al. (2020), TensorFlow Federated (TFF) Google (2019), FederatedScope Kuang et al. (2023); Xie et al. (2023) and Flower Beutel et al. (2020), have adopted recent advancements in FL research. The frameworks especially allow for PEFT of FMs with LoRA. Substra Galtier and Marini (2019), PySyft Ziller et al. (2021), OpenFL Foley et al. (2022), and IBM FL Ludwig et al. (2020), in their versions as of 2023, focus on training smaller FL tasks with 100K up to a 10M parameters and, therefore, do not provide adapters for FM workloads with more than 100M parameters. Yet, a consistent observation across all frameworks is their lack of efficient communication techniques (e.g., FedOBD). Workloads with FMs will significantly increase communication costs, and the growing use cases involving resource-constrained edge and IoT devices require high efficiency for computation and communication. As such, only those FL frameworks enabling training efficiency are viable choices for working with FMs.